The following information was available concerning the revenues of Caterpillar, Inc. and the consumer price index for urban goods (CPI-U).

Year

Revenues

(in thousands)

Average

CPI-U

Constant Dollar Revenues in 2013 Dollars

2010

$42,588

219

2011

$60,138

226

2012

$65,875

230

2013

$55,656

233

7.4 What was the percentage growth rate in revenue during 2013 *after* adjustment for inflation?

-15.5%

-16.6%

20%

-15.8%
Answer #1

Answer = -16.6%

Calculation:

Revenue in 2012 (a) 65875
Revenue in 2013 (b) 55656
Decrease in Revenue = c = b-a (55656 - 65875) -10219
% decrease in revenue = d = c / a (Decrease in Revenue / Revenue for 2012) -15.51%
Average CPI in 2012 (a) 230
Average CPI in 2013 (b) 233
Increase in CPI = c = b-a (233-230) 3
% increase in CPI = d = c/a (Increase in CPI / Average CPI for 2012) 1.30%
% growth in revenue after adjustment of inflation = (1 + revenue increase %) / (1 + Inflation %) - 1
= (1+(-15.51%)) / (1+1.3%) - 1
= (1+(-15.51/100)) / (1+1.3/100) - 1
= (1-0.1551) / (1.013) - 1
= (0.8449 / 1.013) - 1
= 0.834057 - 1
= -0.16594
= (-0.16594*100)%
-16.594%
-16.6%